FEL Upgrade Project Weekly Brief

March 25-29, 2002

 

 

Highlights:

We are pleased to congratulate one of our FEL industrial partners, AES, Inc. of Princeton NJ, who was awarded two JTO grants this week for support of the joint AES/Jefferson Lab project on development of high current injectors for FELs.

Management:

We completed draft CRADAs and work statements for two industrial collaborations and forwarded the documents to DOE for review. One CRADA is in support of the AES 100 mA injector development project; the second CRADA supports the request by AMAC, International for testing of high power rf windows.

Preparations began for the May 1-3 semiannual review of the FEL Upgrade project by discussions on the preliminary agenda, potential presentation material, and likely tour sites for inspection of hardware.

Meetings were held this week on potential revisions to the lab's Environmental Assessment in order to accommodate proposed building modifications for addition of the Helios synchrotron and the planned 100 kW FEL Upgrade to the FEL Facility.

George Neil presented a colloqium at RPI on results from the recently completed FEL user run.

WBS 4 (Injector):

Worked with W&M to resolve arcing problems which occurred with the support tube implantation. Completed preliminary designs on the shielded bellows between the gun and the lightbox and the shielded bellows/BPM combination between the lightbox and the buncher.

Gun HVPS - Waiting for ceramic cylinder to make new Conditioning resistor. The Running resistor has already been received. The mechanical detailing of the parts for the combined mechanism will be underway next week.

WBS 6 (RF):

Quarter HVPS - HV wiring continues on the HVPS for cavity 3. Progress was made this week on the communications between the 2 PLC's that control and readout the various HVPS parameters. Progress was also reported on 485 communications between IOC's and our type of PLC's. We intend to use a similar 485 link in our installation.

Quarter Klystrons - The Machine Shop was finally given the plumbing parts we had on hand for the klystron carts. The balance is on order. A decision was made to use the 2 & 2 1/2 inch ball valves on the new LCW piping. The welding will start again about April 22nd.

Zone 3 - Complete. Waiting for new cryomodule.

Zone 4 - Complete. Waiting for LCW to make final system test on equipment in position 8.

WBS 8 (Instrumentation):

Efforts to organize and track the I&C groups hardware and documentation based on a hand-held bar-code scanner are coming to fruition. The system has now been functionally integrated into a database. It works as follows:

1. Scan any bar-coded item with the hand-held scanned (or any number of items)

2. Key-in any additional notes to describe the item(s)

3. Park the hand-held scanner in it's cradle and press UPLOAD to the pc

4. Open the resultant text file in any text editor to confirm (and/or edit)

5. Logon to http://laser.jlab.org/data and click "Add new items" from the menu on the left.

6. Cut and paste the text file from the hand-held scanned into the form field listed as Paste Barcodes

7. Press Submit (The form will regenerate with the information you pasted)

8. Review the Current Location, Associated System, Hardware Category & Machine Nomenclature pull-downs (All this can be edited later as well and so it is optional at this point)

9. Press the Submit New Items Information button to finish the procedure.

With the help from the operations staff we have a number of the printed circuit board "stuffing jobs" taken care of. The include the VME 4 channel timing module, and the beam viewer control boards. Many thanks!

The beamviewer locations have been assigned to channels / chassis and has been turned over to software for signal names and screen upgrades. The dust covers for the beamviewers have been designed and ordered, the drawing for the beam viewer test box has been turned over to E-cad.

The location for the ion chamber on the wiggler has been identified and D. Dotson has begun work on the communication upgrade to EPICS. (RS-232 to 485). The fourth LeCroy HV pod for the additional Beam Loss Monitors (BLM) has been procured. The Altera programming development for IOC Interfacing and BLM logic is proceeding. EECAD continues to work on documentation for open projects (MPS Input Module, Digital Timing Module, etc.) The main effort has been to get the prototype of the VME Interface stuffed, programmed and working and to continue developing the schematic for the BLM MPS card

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. They are in production mode on coils, devoting two winders to the task. The third coil was potted.

Return legs are in final machining. Top and bottom slabs are out of the heat treat and at the grinding facility. Field clamps remain in process.

We will know next week if the grinding and machining scheduling will work in concert to complete a magnet by the Review.

Injector Dipoles (DU/DV)

Master Machine is cutting up the slab the start rough machining.

WANG NMR is winding all the GU and GV Coils.

Arc Dipoles (GY, GX, GQ)

We received the best and final offers for the GQ-GX magnets. The contracts to very competent vendors are under internal review and are scheduled to be let later today.

AES has been given the Task order to assess the GQ-GX dipoles at lower energy with shorter pole tips. They are also designing a shim system as part of Revision A for the pole tips to reconfigure the magnets for high energy use.

Wang NMR of CA continues making the GY coil winding mandrel and designing the potting mold for the GY coils.

Bosma Machine has the GY core pieces in final planning of the pole surface.

Quadrupoles

3 inch quad (QX)

Production measurements continued on the remaining QX Quadrupoles. The "skinny" quads that David Douglas wants to assign to the regions between cryomodules have very uniform results indistinguishable from the regular series of magnets.

Trim Quad (QT)

Milhous Control of Virginia reports that a first article will be available next Wednesday. Tom Hiatt is scheduled to inspect the magnet.

Sextupole (SF)

DULY Research has generated a successful magnetic design that does not operate in saturation. We have set up a task order with them to make the drawings, copying the successful construction methods of the dispersion section coils and the quadrupole cores.

Octupole (OT)

Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

The configuration for 15% acceptance in the second arc is in the hands of AES and the task order is modified to a wider configuration in those positions. Design at AES continues, drawing up the remainder of the arc chambers.

The drawings of the regions between cryomodules and at the end of the cryomodules are under review for signing on Monday.

Design continued on the details for the telescope region after the cryomodules.

WBS 10 (Wiggler):

We finished calibration measurements of the dispersion section corrector this week. We now have a formula which allows us to set the corrector so that the field integral is zeroed to the 2-3 G-cm level over the working range of the magnet.

We also measured the field with and without the vacuum chamber in place. The field was enhanced by approximately 1 part in 1000. This is greater than one would expect assuming annealed stainless steel. The larger effect is probably due to the amount of work the material endured while being shaped into a rectangular cross section. The strength of the dispersion section and the wiggler will be corrected for the difference.

WBS 11 (Optics):

The purchase order for OCMMS fabricated components was signed and sent to Procurement. Tests of the integrated PZT actuator on the UHV stepper motors (to be used in the optical cavity and collimator) showed performance consistent with specifications. The same PZT actuators were integrated into the OCMMS test stand and operated as expected. Lacking amplifiers (which are still at the vendor), a feedback circuit and amplifier are being constructed. Due to Easter holidays in Europe, the power supply and amplifiers will be received the week of April 8. While most of our efforts are centered on the active stabilization of the cavity mirror orientation, we have continued work on defining the hardware and software required to do the feedback on the cavity mirror ROC.

Design efforts are centered around completing the optical cavity vacuum chambers. The limit switches were moved to provide easier alignment and servicing. The turning mirror cassettes continue to be refined. Thanks to S. Benson's modeling efforts, we have a good idea of the optical layout for the collimator, and can generate specifications for the design. He was also able to complete calculations to define the outcoupler figure, so we can go out for bid on those components. Modeling efforts by J. Gubeli on the 50kW optical beam dump are proceeding.

After evaluating the proposals for the ultrafast laser, an award was made to Spectra-Physics. This will be JLab's first sub 100 fs light source. In parallel with these efforts, the hardware needed (gratings, optomechanics, and single-mode fiber) to do the first stage of compression of "waste" drive laser light is here, and installation will begin today. As a reminder, this hardware will be used for diagnosing electron beam dynamics and their effect on the FEL. In parallel with this effort is one to design an insertable mirror to route the THz light to a point where it can interact with a synchronized ultrafast laser. Calculations show that relative insensitivity of the wavefront of the THz light to misalignment of this mirror.

On the UV cavity, modeling continues on the cryocooled mirrors. The optical penetrations between the vault and upstairs were checked and added to our system print, so we can best determine how to best integrate the UV optical transport wit the IR transport.

Other activities:

We forgot to mention that last week we gave a tour to Mr. W. Decker, Brashear LP. Brashear built the beam director for the THEL project, and is interested in our program. We also held a conference call with J. Albertine and C. Brau on overall progress of the optics WBS in general, and specifically on the performance of the deformable mirror assembly. As a result of this teleconference, we determined there was no need for an on-site visit before the May review.

 

 

--George/Fred/dmag